# A Bird Song Detector for improving bird identification through Deep Learning: a case study from Doñana

### v1

@@ 5 Discussion
The results demonstrate that incorporating a Bird Song Detector significantly enhances bird identification in audio recordings. By segmenting recordings and isolating relevant audio regions, the detector effectively reduces FNs while keeping FPs to a minimum. This pre-filtering process ensures that classifiers operate on cleaner, more focused segments, reducing the risk of misclassification due to background noise or overlapping sounds.
Most publicly available ecoacoustic algorithms (Kahl, 2021; Google Research, 2024) focus primarily on bird species classification rather than the detection of vocalizations. However, general species classification models often struggle in PAM programs deployed in real-world scenarios, as it is already a known case for BirdNET (Pérez-Granados, 2023; Funosas et al., 2024). Our results confirm that relying solely on BirdNET for binary bird detection leads to high FN and FP rates, which limits its ability to detect bird presence.
@@
Compared to BirdNET at 0.6 confidence, the Bird Song Detector significantly improved detection performance. Specifically, FNs decreased by 67% (from 211 to 70), ensuring that far more bird vocalizations were correctly identified. While this came with a 50% increase in FPs (from 6 to 9), the absolute increase was only three additional FPs. In the test subset containing 100-200 minutes of audio, where the potential for false detections is virtually limitless, an increase of three FPs has a negligible impact on overall performance while higher numbers like 63 would create excessive noise in PAM applications.
@@
The Bird Song Detector significantly improved classification accuracy across all models. Without it, classifiers exhibited higher FP rates and greater misclassification errors, as they processed unfiltered, noisy audio segments. The reduction in Idx Pred/Ann indicates that the Bird Song Detector provided cleaner, more structured data, improving the alignment between predictions and ground truth annotations.
@@
However, despite these improvements, classification performance in this study remained lower than the results reported in the original BirdNET research (Kahl et al., 2021). BirdNET was originally evaluated on clean, single-species recordings, achieving a mean average precision of 0.791, an F0.5 score of 0.414 for annotated soundscapes, and an average correlation of 0.251 with hotspot observations (areas with high species diversity). However, in real-world PAM applications, these conditions do not hold, leading to a notable decline in performance (Kahl et al., 2021; Pérez-Granados, 2023). Our findings align with these observations: species classification accuracy remained moderate even after fine-tuning BirdNET and applying the Bird Song Detector. Background noise, species imbalances in training data, and acoustic similarities among certain species contributed to misclassifications, highlighting the ongoing challenges of automated species identification in PAM programs.
@@
These findings suggest that deep learning models pre-trained on image datasets (e.g., ImageNet) do not generalize well to bioacoustic classification tasks. Unlike images, bird vocalizations are temporal and frequency-dependent, requiring models to extract patterns across time and frequency domains. The poor performance of ResNet50 and MobileNetV2 suggests that models optimized for visual features fail to capture the nuances of bioacoustic signals. Future adaptations for deep learning in this field should explore pretraining on bioacoustic datasets rather than visual datasets like ImageNet or using spectrogram-specific architectures designed for temporal pattern recognition (Ghani et al., 2023).
@@
From the dataset split, it is evident that the dataset suffers from a high degree of class imbalance, with some species having a substantially larger number of training samples than others. The poor performance of the model was likely due to this imbalance, as classifiers tend to be biased toward well-represented classes while performing poorly on underrepresented ones (Johnson and Khoshgoftaar, 2019). BirdNET’s built-in oversampling was supplemented with mixup augmentation, but this was insufficient to counteract imbalance, particularly for species with extremely low sample sizes (e.g., Falco tinnunculus or Milvus migrans).
Prior to this upsampling, 17 species had fewer than 10 instances, and their classification results were generally poor, with most having an F1-score of 0.00. In contrast, only three species had more than 50 instances, and they exhibited relatively higher F1-scores, with Sturnus achieving 0.42 and Emberiza calandra reaching 0.46. Class frequency and recall appear highly correlated. A clear example is Anthus pratensis, which had 42 instances in the test set and achieved a recall of 1.00 but a low precision of 0.16, indicating that the model predicts this class frequently. Conversely, species like Cettia cetti and Galerida cristata have moderate recall values (0.13 and 0.47, respectively), showing some predictive ability but still suffering from low precision. This suggests that the model frequently predicted this species, even when incorrect, due to its prevalence in the training data. Conversely, species like Falco tinnunculus (n=2) or Milvus migrans (n=9) had zero recall and precision, indicating that the model was completely unable to recognize them. The model likely overfits to the more frequently occurring classes while underperforming on rare species. This issue is further exacerbated by the presence of species with very small sample sizes, leading to cases where the model never correctly identifies them.
@@
Additionally, BirdNET’s performance is still influenced by biases in its training data. The public datasets used for BirdNET training (e.g., Xeno-canto (canto Foundation, 2024), Macaulay Library (Macaulay Library, 2024)) are dominated by high-quality focal recordings, which lack overlapping sounds and real-world acoustic complexity. This makes BirdNET less effective in identifying species in dense, multi-species environments, such as Doñana National Park.
Although the conservation challenges of Doñana are multiple and require complex solutions from the different social agents involved, we believe that the practical implications of our research for ecological monitoring and conservation of bird populations are significant. Indeed, the integration of Deep Learning models applied to PAM allows for the cost-effective and scalable monitoring of bird diversity, which is essential for tracking bird-species trends over large protected areas. Managers of Doñana can use information from automatic sound classification of bird species to assess the health status of the different habitats present in the area (as many bird species serve as indicators of environmental changes), effectively combining them with direct census monitoring and other methods. This information helps to design conservation measures trying to mitigate bird species diversity loss. At last, the main aim of the BIRDeep project is to enable a real-time tracking of avian diversity to have accurate information on bird diversity in large, protected areas such as Doñana National Park.
@@ 6 Conclusions
We have demonstrated the feasibility and advantages of integrating a Bird Song Detector with fine-tuned classifiers for detecting and classifying bird vocalizations in the soundscapes of Doñana National Park. The proposed two-stage pipeline enhances both detection and classification accuracy by isolating relevant audio segments before species identification. This approach mitigates many of the challenges faced by traditional classification models, particularly in complex acoustic environments.
@@ Acknowledgements
Appendix A Additional Figures and Tables
@@
Precision Recall F1-score Support
@@
Table 2: Classification report for each class of the Bird Song Detector and a fine-tuned BirdNET approach, which was the best performing among all tested models. The metrics include precision, recall and F1-score for each class, as well as the overall accuracy, macro average, and weighted average.

### v2

@@ 5 Discussion
This study presents the methodological development and evaluation of a two-stage pipeline designed to enhance bird species identification in PAM. The results demonstrate that incorporating a Bird Song Detector significantly enhances bird identification in audio recordings. By segmenting recordings and isolating relevant audio regions, the detector effectively reduces FNs while keeping FPs to a minimum. This pre-filtering process ensures that classifiers operate on cleaner, more focused segments, reducing the risk of misclassification due to background noise or overlapping sounds.
Evaluation of Detectors. Most publicly available ecoacoustic algorithms (Kahl, 2021; Google Research, 2024) focus primarily on bird species classification rather than the detection of vocalizations. However, general species classification models often struggle in PAM programs deployed in real-world scenarios, as it is already a known case for BirdNET (Pérez-Granados, 2023; Funosas et al., 2024). Our results confirm that relying solely on BirdNET for binary bird detection leads to high FN and FP rates, which limits its ability to detect bird presence.
Selecting an appropriate confidence threshold for the Bird Song Detector is key to balance recall and precision. As shown, increasing the threshold reduces FPs but also results in substantial TP losses—up to 99% at the highest setting. We found that a threshold of 0.15 offered the best compromise, maintaining reasonable recall while limiting the number of non-informative detections processed downstream. This balance is critical in real-world deployments, where excessively low thresholds would flood subsequent stages with noise, while overly conservative thresholds risk missing important vocalizations. The chosen threshold of 0.15 minimizes FN rates without overwhelming the classifier with FP detections, thus optimizing overall system efficiency and ecological reliability.
@@
Threshold Optimization and Error Analysis. Compared to BirdNET at 0.6 confidence, the Bird Song Detector significantly improved detection performance. Specifically, FNs decreased by 67% (from 211 to 70), ensuring that far more bird vocalizations were correctly identified. While this came with a 50% increase in FPs (from 6 to 9), the absolute increase was only three additional FPs. Considering that this evaluation was conducted on the test subset, which comprised approximately 100 to 200 minutes of annotated audio, where the potential for false detections is virtually limitless, an increase of three FPs has a negligible impact on overall performance while higher numbers like 63 would create excessive noise in PAM applications.
@@
Comparison of Classifiers and Deep Learning Models. The Bird Song Detector significantly improved classification accuracy across all models. Without it, classifiers exhibited higher FP rates and greater misclassification errors, as they processed unfiltered, noisy audio segments. The reduction in Idx Pred/Ann indicates that the Bird Song Detector provided cleaner, more structured data, improving the alignment between predictions and ground truth annotations.
@@
However, despite these improvements, classification performance in this study remained lower than the results reported in the original BirdNET research (Kahl et al., 2021). BirdNET was originally evaluated on clean, single-species recordings, achieving a mean average precision of 0.791, an F0.5 score of 0.414 for annotated soundscapes, and an average correlation of 0.251 with hotspot observations (areas with high species diversity). However, in real-world PAM applications, these conditions do not hold, leading to a notable decline in performance (Kahl et al., 2021; Pérez-Granados, 2023). Our findings align with these observations, species classification accuracy remained moderate even after fine-tuning BirdNET and applying the Bird Song Detector. Background noise, species imbalances in training data, and acoustic similarities among certain species contributed to misclassifications, highlighting the ongoing challenges of automated species identification in PAM programs.
@@
These findings suggest that deep learning models pre-trained on image datasets (e.g., ImageNet) do not generalize well to bioacoustic classification tasks. Unlike images, bird vocalizations are temporal and frequency-dependent, requiring models to extract patterns across time and frequency domains. The poor performance of ResNet50 and MobileNetV2 suggests that models optimized for visual features fail to capture the nuances of bioacoustic signals. Future adaptations for deep learning in this field should explore pretraining on bioacoustic datasets rather than visual datasets like ImageNet or using spectrogram-specific architectures designed for temporal pattern recognition (Ghani et al., 2023; Xiao et al., 2022; Xie et al., 2022; Gong et al., 2021).
@@
Limitations Due to Class Imbalance. One limitation of our dataset is the imbalance in the number of annotated vocalizations per species. This imbalance mirrors natural patterns in animal abundance, behavior, and detectability — a well-documented issue in ecological monitoring studies (Cui et al., 2019; Beery et al., 2018; MacKenzie et al., 2002). While this poses a challenge for classification performance, it also reflects the operational conditions of real-world PAM applications, reinforcing the need for methods that can handle such skewed data distributions.
This class imbalance was also evident during dataset splitting and model evaluation. Some species had a disproportionately larger number of training samples than others, while others, such as Falco tinnunculus and Milvus migrans, were extremely underrepresented. The poor performance of the model was likely due to this imbalance, as classifiers tend to overfit to these dominant classes while performing poorly on underrepresented ones (Johnson and Khoshgoftaar, 2019; Cui et al., 2019; Pantazis et al., 2024). Although BirdNET’s built-in oversampling was supplemented with mixup augmentation, these techniques were not sufficient to counteract imbalance effect—particularly for underrepresented species. Combining diverse augmentation methods has shown promising results in recent work for improving species-level recall in imbalanced datasets (Kumar et al., 2024).
Prior to this upsampling, 17 species had fewer than 10 instances, and their classification results were generally poor, with most having an F1-score of 0.00. In contrast, only three species had more than 50 instances, and they exhibited relatively higher F1-scores, with Sturnus achieving 0.42 and Emberiza calandra reaching 0.46. Class frequency and recall appear highly correlated. A clear example is Anthus pratensis, which had 42 instances in the test set and achieved a recall of 1.00 but a low precision of 0.16, indicating that the model predicts this class frequently. Conversely, species like Cettia cetti and Galerida cristata have moderate recall values (0.13 and 0.47, respectively), showing some predictive ability but still suffering from low precision. This suggests that the model frequently predicted this species, even when incorrect, due to its prevalence in the training data. Conversely, species like Falco tinnunculus (n=2) or Milvus migrans (n=9) had zero recall and precision, indicating that the model was completely unable to recognize them. These bird species normally fly over large altitude and far away from the recorders, so getting more vocalizations that could be labelled and used to train the models would require a more focused sampling. The model likely overfits to the more frequently occurring classes while underperforms on rare species. This issue is further exacerbated by the presence of species with very small sample sizes, leading to cases where the model never correctly identifies them.
@@
Data Bias in Public Bird Sound Resources. Additionally, BirdNET’s performance is still influenced by biases in its training data. The public datasets used for BirdNET training (e.g., Xeno-canto (canto Foundation, 2024), Macaulay Library (Macaulay Library, 2024)) are dominated by high-quality focal recordings, which lack overlapping sounds and real-world acoustic complexity. This makes BirdNET less effective in identifying species in dense, multi-species environments, such as Doñana National Park.
Ecological Implications and Future Applications. Although the conservation challenges of Doñana are multiple and require complex solutions from the different social agents involved, we believe that the practical implications of our research for ecological monitoring and conservation of bird populations are significant. Indeed, the integration of Deep Learning models applied to PAM allows for the cost-effective and scalable monitoring of bird diversity, which is essential for tracking bird-species trends over large protected areas. Managers of Doñana can use information from automatic sound classification of bird species to assess the health status of the different habitats present in the area (as many bird species serve as indicators of environmental changes), effectively combining them with direct census monitoring and other methods. This information helps to design conservation measures trying to mitigate bird species diversity loss. While the current work is based on a subset of annotated data, full-scale deployment of this pipeline across the entire dataset from Doñana National Park is ongoing. For this reason, winter migratory species may not be represented in the annotated subset used here. However, future analyses will incorporate these additional recordings. At last, the main aim of the BIRDeep project is to enable a real-time tracking of avian diversity to have accurate information on bird diversity in large, protected areas such as Doñana National Park.
@@ 6 Conclusions
We have demonstrated the feasibility and advantages of integrating a Bird Song Detector with fine-tuned classifiers for detecting and classifying bird vocalizations in the soundscapes of Doñana National Park. The proposed pipeline enhances both detection and classification accuracy by isolating relevant audio segments before species identification. This approach mitigates many of the challenges faced by traditional classification models, particularly in complex acoustic environments.
@@ Acknowledgements
Appendix A Additional Technical Details
YOLOv8 Annotation Format and File Structure
Annotations were created by manually labeling the temporal and frequency localization of vocalizations—i.e., using start and end times (in seconds) and frequency ranges (in Hz) relative to each 60-second audio file. These real-world values were then transformed into bounding boxes formatted according to the YOLOv8 object detection standard, where all coordinates are normalized with respect to the spectrogram image dimensions.
  • 1.
    class: the zero-indexed label corresponding to the annotated class,
  • 2.
    x_center, y_center: the normalized center coordinates of the bounding box,
  • 3.
    x_width, y_height: the normalized width and height of the bounding box.
Due to data limitations and to simplify the detection task, we chose to fix the vertical dimension of all bounding boxes to span the full frequency range of the spectrogram. This means each bounding box covers the entire vertical axis, and only the horizontal (temporal) dimension is variable. Thus, the vertical values were set as:
$y\_center=0.5,\quad y\_height=1.0$
Annotations were stored in individual .txt files, each corresponding to a spectrogram image. These files were placed in a mirrored directory structure matching the organization of the original audio files. In accordance with YOLOv8 standards (Ultralytics, 2024), each file contains one row per annotation in the following format:
$\textit{class}\quad x\_center\quad y\_center\quad x\_width\quad y\_height$
Spectrograms with no annotated vocalizations have an empty .txt file.
Temporal Conversion of Bird Song Detector Predictions
Since the Bird Song Detector is implemented as a YOLOv8 object detector, its output consists of predicted bounding boxes in YOLOv8 format. To interpret these predictions in ecologically meaningful terms—i.e., as time intervals in the original audio—they must be converted back from normalized image coordinates to temporal coordinates in seconds.
Each spectrogram image corresponds to a 60-second audio segment and has fixed dimensions of $W=930$ pixels in width (time axis) and $H=462$ pixels in height (frequency axis). Therefore, the temporal resolution is approximately 0.0645 seconds per pixel.
YOLOv8 bounding boxes are returned in normalized format. Each box provides:
  • 1.
    $x_{\text{center}}$: the normalized center of the box along the time axis,
  • 2.
    $w$: the normalized width of the box.
To compute the corresponding temporal boundaries in seconds, the following denormalization and conversion formulas are used:
$x_{\textit{center\_d}}=x_{\textit{center}}\times\textit{W}$
[…]
Where $W$ is the width of the spectrogram image used as input for the Bird Song Detector, set to $W=930$ pixels in our case. The variables $x_{\text{center}}$ and $w$ are normalized values representing the bounding box center and width, respectively. To map these predictions to real-world time, they are first denormalized into $x_{\textit{center\_d}}$ and $w_{\textit{d}}$ by scaling with $W$. The resulting values, start_sec and end_sec, correspond to the start and end times (in seconds) of the detected bird song segment relative to that segment.
As shown in Figure 1, bounding boxes are first normalized for training and then later converted back into real-world time using the spectrogram width and duration. This visualization helps clarify the role of each variable in the forward and reverse transformation, and illustrates how the detector’s output aligns with the original audio.
Figure 1: Visual representation of the YOLOv8 annotation format and its conversion between image-based bounding boxes and real-time bird vocalization intervals. The 60-second spectrogram shown includes an example annotation of a vocalization spanning from 30.0 to 35.0 seconds, corresponding to a width of 930 pixels (W). This annotated vocalization is normalized and converted to temporal coordinates based on YOLOv8 outputs, using the formulas displayed on the left side of the image. Conversely, YOLOv8 bounding boxes can be reconverted to temporal annotations using the formulas in the center of the image. Since all vocalizations are annotated across the full frequency range, the vertical axis of the bounding box remains constant across samples: $y_{\text{center}}=0.5$, representing the normalized center of the image, and $y_{\text{height}}=1.0$, spanning the full height of the image.
Appendix B Additional Figures and Tables
@@
Precision Recall F1-score Number of test samples
@@
Table 2: Classification report for each class of the Bird Song Detector and a fine-tuned BirdNET approach, which was the best performing among all tested models. The table presents precision, recall and F1-score for each class. The last three rows show overall performance metrics: Accuracy, Macro average, and Weighted average.